OpenStudy (dls):

x= a (theta - sin theta) y=a(1-cos theta) dy/dx=?

OpenStudy (dls):

my attempt: \[\LARGE \frac{dy}{dx}=\frac{\sin \theta}{1-\cos \theta}\]

OpenStudy (dls):

On differentiating x I got : a(1-cos theta) On differentiating y I got : (asin theta)

hartnn (hartnn):

seems correct....

OpenStudy (dls):

WOAH :O thanks!

hartnn (hartnn):

u want in terms of x and y ?

OpenStudy (dls):

yes I guess

OpenStudy (dls):

I just have to find dy/dx idk

hartnn (hartnn):

then leave it..